# Shapeng Jiang Lijia Wei Chen Zhang

### v1

@@
Donald Trumps in the Virtual Polls: Simulating and Predicting Public Opinions in Surveys Using Large Language Modelsthanks: All authors have contributed equally and are listed in alphabetical order by last name. Corresponding author: Lijia Wei, ljwei@whu.edu.cn
@@
In recent years, large language models (LLMs) have attracted attention due to their ability to generate human-like text. As surveys and opinion polls remain key tools for gauging public attitudes, there is increasing interest in assessing whether LLMs can accurately replicate human responses. This study examines the potential of LLMs, specifically ChatGPT-4o, to replicate human responses in large-scale surveys and to predict election outcomes based on demographic data. Employing data from the World Values Survey (WVS) and the American National Election Studies (ANES), we assess the LLM’s performance in two key tasks: simulating human responses and forecasting U.S. election results. In simulations, the LLM was tasked with generating synthetic responses for various socio-cultural and trust-related questions, demonstrating notable alignment with human response patterns across U.S.-China samples, though with some limitations on value-sensitive topics. In prediction tasks, the LLM was used to simulate voting behavior in past U.S. elections and predict the 2024 election outcome. Our findings show that the LLM replicates cultural differences effectively, exhibits in-sample predictive validity, and provides plausible out-of-sample forecasts, suggesting potential as a cost-effective supplement for survey-based research.
@@ 1 Introduction
Large language models (LLMs) are powerful tools for simulating human behavior due to their ability to capture the complexity of natural language and encode a wide range of human experiences, cultural norms, and decision-making patterns from extensive training data. These models reflect how people use language to express thoughts, beliefs, and emotions, making them suitable for replicating behaviors across different social contexts. This is particularly valuable in social science research, where experiments and surveys rely heavily on human participants to gather behavioral data (Kim and Lee,, 2023). Specifically, using LLMs in behavioral simulations allows researchers to scale experiments, lower the costs of human subject studies, and gain insights into individual behavior and complex social interactions, such as trust, negotiation, and cooperation (Aher et al.,, 2022). Additionally, LLMs can simulate multiple scenarios and personalities, helping researchers predict responses to new social environments (Cheng et al.,, 2023).
@@
LLMs are primarily trained on vast text corpora derived from diverse sources, such as general web text, books, and articles, rather than raw experimental data. Despite this, current research on LLM-based simulations found that LLMs can accurately reproduce outcomes from classic experiments such as the ultimatum game, trust game, prisoner’s dilemma, dictator game, and strategic communication game (Aher et al.,, 2022; Xie et al.,, 2024; Phelps and Russell,, 2023; Xu et al.,, 2023; Ashokkumar et al.,, 2024). However, current LLM-based experimental simulations focus on basic experiments that are relatively simple, with clear incentives or dominant strategies. Compared to experimental simulations, surveys feature ambiguously defined questions with multiple options that offer minimal distinction (e.g., Likert scales). Without incentives and optimal strategies, surveys focus more on values and attitudes, where consensus may not exist.
This study aims to leverage the latest LLM model, ChatGPT-4o [^1] The ChatGPT-4o model has been trained on information available until October 2023. Throughout the entire study, we utilized the ChatGPT API tool exclusively, without incorporating any real-time search capabilities., to simulate respondents’ choices in large-scale surveys, particularly their attitudes on controversial value-based questions. Specifically, we leveraged the LLM to perform two main tasks: simulation and prediction. For the simulation task, the model adopted personas based on demographic profiles from respondents in the U.S. and China, using data from the World Values Survey (WVS). Based on the adopted personas in the WVS, the model generated synthetic responses on various topics such as social values, trust, and ethics, allowing us to compare the synthetic responses with actual human responses. In contrast, for the prediction task, the LLM simulated voting behavior using another group of personas sourced from the American National Election Studies (ANES), first by recreating past U.S. election outcomes and then forecasting the 2024 election.
Our findings reveal that the LLM shows strong in-sample predictive ability, closely aligning with human responses when demographic inputs are included. For cross-cultural comparisons, the LLM effectively replicates cultural differences in values between the U.S. and China on most items, although it occasionally overestimates socially progressive views, particularly in the U.S. On the predictive side, the model demonstrates potential for out-of-sample forecasting, with its accuracy further supported when historical voting patterns are incorporated. These results underscore the potential of LLMs as supplemental tools in large-scale surveys, enabling cost-effective and culturally nuanced insights.
@@
Literature have demonstrated the ability of LLMs in simulating respondents’ choices in surveys. Studies such as Argyle et al., (2023) explored how GPT-3, conditioned with socio-demographic data, can generate responses resembling real-world survey data, capturing nuanced relationships between attitudes and socio-cultural contexts . Similarly, Bisbee et al., (2023) evaluated LLM’s ability to generate synthetic opinions aligned with real survey responses, finding that while overall averages can match human data, the model struggles with variability and reliability across iterations. Research also highlights the practical challenges of using LLMs in surveys; for example, Kim and Lee, (2023) noted that LLM-generated survey responses may exhibit response biases based on question phrasing, which can affect the reliability of insights. Furthermore, Tjuatja et al., (2023) emphasized that while LLMs can replicate survey patterns, they tend to homogenize responses, reducing the distinctiveness found in human data. These findings illustrate both the promise and limitations of using LLMs to simulate survey participants, underscoring the need for careful prompt design and attention to variability. However, current research on LLM-based survey simulations is still limited, with most studies focused on political attitudes and data from English-speaking countries.
Our paper contributes to the literature by shifting the focus of LLM-based simulations from experimental replication to survey studies. We employ the more advanced ChatGPT-4o to explore key social values and opinions, broadening the relevance to social science surveys. Our study also offers cross-cultural insights through U.S.-China comparisons, providing rare evidence from non-English-speaking contexts. Additionally, we leverage simulated results to predict the 2024 U.S. presidential election based on voters’ demographics—marking the first use of LLMs to forecast a future event.
2 Research design
This section outlines the research design for executing the two tasks: simulation and prediction. In the simulation task, we provided ChatGPT with demographic characteristics of respondents from the United States and China, as recorded in the 7th Wave of the World Values Survey (WVS). After adopting a persona defined by a set of demographic characteristics of a respondent, the LLM was tasked with generating a dataset of synthetic opinions on social values, trust, common-sense questions, and ethical norms and values. By doing so, we can compare the synthetic opinions with the true opinions provided by human respondents in the survey. This not only enables us to assess the accuracy of the synthetic LLM responses but also to determine whether the opinion differences between China and the U.S. are retained in the simulated sample. The prompts used in this task is presented in the part (a) of Appendix A. In total, we compare the LLM’s responses with those of each of the 2,429 human respondents in China and 2,507 in the U.S. from the survey.
In the prediction task, the LLM was provided with the demographic characteristics of respondents from the American National Election Studies (ANES). We prompted it to play the role of an ANES respondent and cast a vote accordingly. After adopting the persona of a respondent from the 2020 ANES survey, the LLM was tasked with voting in two elections: Hillary Clinton vs. Donald Trump in 2016 and Joe Biden vs. Donald Trump in 2020. Following U.S. election rules, we then aggregated the simulated votes by state. These were further combined at the national level to determine the electoral votes and predict the winner. Thus, we can compare the synthetic voting results generated by the LLM with the actual outcomes of the elections. Beyond that, we prompted the LLM to adopt the persona of a respondent from the 2020 ANES survey and cast a hypothetical vote in the 2024 presidential election, choosing between Kamala Harris and Donald Trump. Therefore, after the official election results are released, we can verify the accuracy of our predictions on a state-by-state basis. In total, we utilized the personal characteristics of 6,571 human respondents from the 2020 ANES survey and adjusted the prediction results using the survey’s sampling weights to ensure representativeness.
Since the LLM is highly sensitive to prompts, we employed two different prompting methods. The first prompt is called the “role-play prompt”, a widely used approach in LLMs. It involves assigning the model a specific role and instructing it to act based on the characteristics and information associated with that role. This technique helps guide the model’s behavior and responses, making them more contextually appropriate and aligned with the task at hand. As shown in part (b) of Appendix A, we instruct the LLM to play the role of a resident being interviewed by a pollster.
However, using a role-play prompt for predicting voting behavior may lead the LLM to focus too heavily on individual characteristics while neglecting the historical voting trends and established political leanings of a given state. To address this, we implemented a “structural prompt,” guiding the LLM to consider both the respondent’s profile and the political tendencies of their state. This prompt provides more structured background information and directs GPT to incorporate the state’s historical voting patterns into its predictions. For the detailed prompt, see part (c) of Appendix A.
3 Results
3.1 LLM accuracy in replicating U.S. and China survey responses
@@ 3 Results
Figure 1 compares human responses from the WVS sample with those generated by the LLM. Throughout this section, survey questions are referenced directly by their original identifiers in the WVS. Panel A displays the U.S. sample, and Panel B shows the China sample. While there are some mean differences between LLM and human responses, most LLM response means fall within one standard deviation of the human response means. Additionally, many LLM responses not only have means within this range but also exhibit a smaller variance, as their own standard deviation is largely encompassed by that of the human responses, indicating less fluctuation in LLM responses. This finding is consistent with Bisbee et al., (2023).
Panel A: WVS-U.S.
Panel B: WVS-China
@@
However, in some survey items, the LLM responses diverge notably from human responses. For instance, in the U.S. sample, significant differences appear on social values questions such as Q33, Q36, and Q41, which address topics like gender equality in employment, LGBTQ+ rights, and work ethic. For trust-related items (Q62 and Q63), the LLM predicts a higher level of trust toward individuals from diverse religious and national backgrounds than actual U.S. respondents expressed. These findings reflect that ChatGPT may exhibit a bias toward portraying Americans as more inclined toward progressive views, aligning with observations from Feng et al., (2023). Interestingly, for these questions, the differences between LLM and human responses are smaller in the China sample, indicating the LLM’s stronger alignment with values in Chinese context and a less noticeable emphasis on socially progressive views.
For common knowledge items like Q91 and Q92, the LLM generally outperforms human respondents with minimal variation, likely due to its vast training data. The LLM also aligned well with human responses on ethical norms, and it tends to show less variability on moral questions. An exception is China’s Q182, where the LLM notably overestimates societal acceptance of homosexual communities, reflecting a possible data-influenced overprojection of acceptance.
Table 1 shows differences between U.S. and China responses in both human and LLM samples. We observe that most cultural differences found in the human data persist in the LLM-generated responses, maintaining the same directional trends with statistically significant $t$-test results. In Panel A, the LLM replicates all actual U.S.-China differences in social values. In Panel B, Q60 and Q61 diverge in predicted direction. Panel C, focusing on common-sense knowledge, shows LLM responses are stable and unrelated to U.S.-China differences. In Panel D, the LLM inconsistently captures U.S.-China differences, retaining them only for Q182, Q185, and Q190. The weaker performance in Panel D may stem from the LLM’s tendency to align with established laws and social norms, which can cause it to diverge from human responses.
Table 1: U.S.-China differences: human responses vs. LLM responses
Human response LLM response
U.S. China Difference U.S. China Difference
@@
Q33 3.764 3.022 0.742 4.828 2.982 1.846
Q34 2.913 2.247 0.666 3.045 1.781 1.264
Q35 3.628 3.414 0.214 4.242 3.544 0.698
Q36 2.351 3.653 -1.302 1.292 3.252 -1.96
Q37 3.672 2.272 1.4 3.806 1.797 2.009
Q38 2.746 1.405 1.341 2.224 0.99 1.234
Q39 2.723 1.968 0.755 3.357 2.313 1.044
Q40 2.434 2.028 0.406 4.864 1.667 3.197
Q41 3.28 2.12 1.16 4.392 1.873 2.519
@@
Q58 1.443 1.147 0.296 1.04 1 0.04
Q59 2.222 2.022 0.2 2.028 1.996 0.032
Q60 1.788 1.996 -0.208 1.966 1.906 0.06
Q61 2.741 3.044 -0.303 3.023 2.997 0.026
Q62 2.216 3.028 -0.812 2.931 2.994 -0.063
Q63 2.202 3.084 -0.882 2.915 3.106 -0.191
Panel C: Common-sense on international organization
Q91 1.901 2.329 -0.428 3 2.999 0.001
Q92 1.513 1.57 -0.057 1 0.999 0.001
Panel D: Ethical norms and values
Q178 2.843 1.62 1.223 1.421 1.824 -0.403
Q179 1.847 1.283 0.564 1 1 0
Q180 2.059 1.5 0.559 1.29 1.875 -0.585
Q181 1.714 1.595 0.119 1 1.072 -0.072
Q182 6.552 2.331 4.221 8.031 5.676 2.355
Q185 6.633 3.738 2.895 6.93 5.442 1.488
Q190 1.926 3.325 -1.399 1.005 2.375 -1.37
Q192 1.51 1.317 0.193 1 1 0
Note: * denote statistical significance at the 1% level.
3.2 Variance-covariance of responses and similarity of correlations
@@
Researchers often care about correlations between behaviors and opinions. Due to budget constraints, they often measure multiple preferences in one survey or expriment (e.g., trust and risk-taking), making the consistency of pairwise correlations among these measurements important. After evaluating how well the LLM matches human responses when responding to individual questions, we also explored whether it retains the correlations between different responses observed in the human sample.
We use a method similar to that of Snowberg and Yariv, (2021) to analyze the relationships between responses to different questions in both LLM-simulated and human samples. Specifically, we assess whether statistically significant pairwise correlations observed between two questions in the LLM-generated responses are likewise present in human responses. The results, shown in Figure 2, display the direction and significance of these pairwise correlations for questions on social values and trust, with panel (a) representing the U.S. sample and panel (b) the Chinese sample.
@@
A positive ($+$) or negative (--) and significant correlation between two outcome measures is denoted by “$+$” or “--”, while an insignificant correlation is marked with “0”. Each cell contains two signs: the first indicates correlations in the LLM sample, and the second represents those in the human sample. Cells where pairwise correlations agree in both sign and significance between LLM-simulated and human samples are labeled as “complete agreement” and shaded light grey. Cells where the correlation is significant in one sample but insignificant in the other are labeled “partial disagreement” and shaded grey. Lastly, cells with significant correlations in both samples but in opposite directions are marked as “complete disagreement” and shaded dark grey. We do not include common-sense ethical norms and values questions because LLM-simulated responses lack variation.
(a) Social values
(b) Trust
Panel A: WVS-U.S.
(c) Social values
(d) Trust
Panel B: WVS-China

### v2

@@
Donald Trumps in the Virtual Polls: Predicting Public Opinions Using Large Language Modelsthanks: All authors have contributed equally and are listed in alphabetical order by last name. Corresponding author: Lijia Wei, ljwei@whu.edu.cn
@@
In recent years, large language models (LLMs) have attracted attention due to their ability to generate human-like text. As surveys and opinion polls remain key tools for gauging public attitudes, there is increasing interest in assessing whether LLMs can accurately replicate human responses. This study examines the potential of LLMs, specifically ChatGPT-4o, to replicate human responses in large-scale surveys and to predict election outcomes based on demographic data. Employing data from the World Values Survey (WVS) and the American National Election Studies (ANES), we assess the LLM’s performance in two key tasks: predicting human survey responses and U.S. election results. In survey tasks, the LLM was tasked with generating synthetic responses for various socio-cultural and trust-related questions, demonstrating notable alignment with human response patterns across U.S.-China samples, though with some limitations on value-sensitive topics. In voting tasks, the LLM was mainly used to simulate voting behavior in past U.S. elections and predict the 2024 election outcome. Our findings show that the LLM replicates cultural differences effectively, exhibits in-sample predictive validity, and provides plausible out-of-sample forecasts, suggesting potential as a cost-effective supplement for survey-based research.
@@
Large language models (LLMs) are powerful tools for predicting human behavior due to their ability to capture the complexity of natural language and encode a wide range of human experiences, cultural norms, and decision-making patterns from extensive training data. These models reflect how people use language to express thoughts, beliefs, and emotions, making them suitable for replicating behaviors across different social contexts. This is particularly valuable in social science research, where experiments and surveys rely heavily on human participants to gather behavioral data (Kim and Lee,, 2023). Specifically, using LLMs in behavioral predictions allows researchers to scale experiments, lower the costs of human subject studies, and gain insights into individual behavior and complex social interactions, such as trust, negotiation, and cooperation (Aher et al.,, 2022). Additionally, LLMs can simulate multiple scenarios and personalities, helping researchers predict responses to new social environments (Cheng et al.,, 2023).
Previous literature has primarily focused on using large language models (LLMs) for prediction tasks, relying on training data derived from “rational” materials such as general web text, books, and articles. However, the current development trajectory of LLMs is increasingly emphasizing reasoning and rationality(Qi et al.,, 2024; Zhang et al.,, 2024), which fundamentally differs from the core requirements of public opinion prediction. Public opinion prediction inherently involves non-rational factors, such as social influences, cognitive biases, and emotional responses, which cannot be adequately captured by models trained solely on existing texts and similar materials.  [^1] As LLMs continue to evolve towards more “rational” models, the challenge remains that the presence of these non-rational elements in group decision-making will inevitably introduce errors in predictions, making it impossible to fully resolve the discrepancies between model outputs and actual public opinions.
In his book Thinking, Fast and Slow, Daniel Kahneman describes human thinking as reliant on two distinct systems: System 1, which is fast, intuitive, and influenced by tradition and experience, and System 2, which is slow, rational, and analytical (Kahneman,, 2011). This duality in human cognition is crucial for understanding the decision-making processes of individuals, particularly when they express social values or cast votes. In a similar vein, large language models (LLMs) are increasingly designed to simulate the rational, slow thinking of humans, processing vast datasets and generating responses based on reasoning and logic. However, this focus on rationality contrasts with the inherently non-rational, fast thinking that influences public opinions.
Our study introduces the Matching-LLM method [^2] The term “Matching” in the Matching-LLM method carries a dual meaning. First, it involves pairing each real human respondent with an LLM-generated counterpart. Second, it matches each real human respondent with a historical respondent from past survey data., which seeks to bridge this gap by combining the “System 2” of LLMs with the “System 1” shaped by human traditions and intuition, derived from historical data. By integrating both types of cognitive processing, we aim to more accurately model the complexities of collective human behavior, particularly in predicting public opinions. Additionally, we place a strong emphasis on incorporating historical and cultural factors into the synthesis of survey responses with LLMs. This approach acknowledges that respondents’ opinions are influenced not only by slow, rational thinking but also by the subconscious, fast thinking shaped by their regional and cultural context (Nisbett et al.,, 2001; Lodge,, 2013).
Our study is comprised of two main tasks: the survey prediction and the election prediction. For the survey task, the model adopted personas based on demographic profiles from respondents in the U.S. and China, using data from the World Values Survey (WVS). The WVS was chosen because it covers important topics related to social values and includes respondents from multiple countries, enabling cross-national comparisons. Based on these personas, the model generated synthetic responses on topics such as gender role, work ethics, trust, and morality, allowing us to compare the synthetic responses with actual human responses. For the voting task, we study the performance of LLM on predicting electoral voting. First, the model used personas from the American National Election Studies (ANES), a dataset renowned for detailed voter demographics, to recreate the outcomes of the 2016 and 2020 U.S. elections. Second, based on the voter characteristics of the 2020 U.S. presidential election, the LLM was asked to generate votes for the 2024 presidential election between Donald Trump and Kamala Harris. This prediction was completed prior to the election, and after the election results were announced, we compared the actual outcomes with the LLM’s predictions. This comparison allowed us to evaluate the model’s out-of-sample predictive capability.
To preview our results, the Matching-LLM method significantly outperforms traditional LLM approaches. The Matching-LLM effectively simulates people’s opinions on social values and captures the correlations among these questions, as well as cross-cultural differences. Additionally, the Matching-LLM voting behavior not only replicated past U.S. election outcomes but also predicted Donald Trump’s overwhelming victory in the 2024 U.S. presidential election, closely aligning with the actual results across most states. The incorporation of historical factors, which reflect human traditions and intuition derived from historical data, significantly enhanced the LLM’s performance in both tasks.
@@
Current research on LLM-based predictions found that LLMs can accurately reproduce outcomes from classic experiments such as the ultimatum game, trust game, prisoner’s dilemma, dictator game, and strategic communication game (Aher et al.,, 2022; Xie et al.,, 2024; Phelps and Russell,, 2023; Xu et al.,, 2023; Ashokkumar et al.,, 2024). However, current LLM-based experimental predictions focus on basic experiments that are relatively simple, with clear incentives or dominant strategies. Compared to experimental predictions, surveys usually feature ambiguously defined questions with multiple options that offer minimal distinction (e.g., Likert scales). Without incentives and optimal strategies, surveys focus more on values and attitudes, where consensus may not exist.
@@
Literature have demonstrated the ability of LLMs in predicting respondents’ choices in surveys. For instance, GPT-3, an early version of a LLM trained on data up to mid-2020, can generate responses resembling real-world survey data and capture nuanced relationships between attitudes and socio-cultural contexts (Argyle et al.,, 2023). Research also highlights the practical challenges of using LLMs in surveys. Specifically, while LLM-generated survey responses can align with real survey averages, the LLMs often struggle with variability, exhibit biases influenced by question phrasing, and homogenize responses, thereby reducing the distinctiveness observed in human data Bisbee et al., (2023); Kim and Lee, (2023); Tjuatja et al., (2023). These findings illustrate both the promise and limitations of using LLMs to simulate survey participants, underscoring the need for careful prompt design and attention to variability. In addition, these studies primarily focus on political attitudes rather than social values and are limited to predicting survey responses from English-speaking countries.
Results
WVS survey responses: in-sample prediction
Mean comparisons. Figure 1 compares human responses from the WVS sample with those generated by the LLM. Throughout this section, survey questions are referenced using concise phrases as summaries of the original questions. The figure illustrates three scenarios represented by different colors: red indicates the actual responses from human samples in WVS7, black represents the LLM-generated responses based on the demographic characteristics of the human samples in WVS7, and green depicts the weighted results, referred to as Matching-LLM, combining the LLM-generated responses (black) with historical response data from WVS6. We applied 1:1 nearest neighbor matching in the Propensity Score Matching (PSM) method to identify the most similar respondent from WVS6 for each respondent in WVS7. The responses of the matched pairs were then weighted to generate the Matching-LLM result [^3] Since Matching-LLM relies on sample information from WVS6, the results are only shown for questions common to both WVS6 and WVS7. Most of the survey questions are 3-to-10-point Likert scales..
@@ Results
While there are some mean differences between LLM and human responses, most LLM-synthesized or Matching-LLM-synthesized response means fall within one standard deviation of the human response means. In addition, many LLM or Matching-LLM responses exhibit a smaller variance, as their own standard deviation is largely encompassed by that of the human responses, indicating less fluctuation in LLM responses. This finding is consistent with previous research on synthetic responses(Bisbee et al.,, 2023).
@@
However, in some survey items, the LLM responses diverge notably from human responses. For instance, in the U.S. sample, significant differences appear on social values questions such as “Homosexual parents”, and “Work over leisure”, which address topics like LGBTQ+ rights, and work ethic. For trust-related items (“Trust on religious outgroups” and “Trust on foreigners”), the LLM predicts a higher level of trust toward individuals from diverse religious and foreign countries than actual U.S. respondents expressed. These findings reflect that ChatGPT may exhibit a bias toward portraying Americans as more inclined toward progressive views, aligning with observations from prior studies on LLM-synthesized responses(Feng et al.,, 2023). Interestingly, for these questions, the differences between LLM and human responses are smaller in the China sample, indicating the LLM’s alignment with Chinese values and a less noticeable emphasis on socially progressive views. An exception is “Homosexual acceptance,” where the LLM still exhibits a high level of acceptance that does not align with the Chinese context.
For common knowledge items like “Question on the IMF” and “Question on the UN”, the LLM generally outperforms human respondents with minimal variation, due to its vast training data. The LLM-generated responses also tend to show less variability on moral or law questions. For example, the LLM show no tolerance for theft, bribery, and terrorism. Interestingly, when the LLM played the role of U.S. respondents, it clearly rejected parental abuse with barely any variation, yet when predicting Chinese respondents, it showed a certain level of acceptance. This result may stem from differences in parenting styles between China and the United States(Chao,, 1994), while the lack of variation in U.S. responses may be attributed to adherence to U.S. laws.
Historical Data-enhanced LLM. As shown in Figure 1, the response means of the Matching-LLM generally align more closely with the human response means in WVS7 compared to the raw LLM. However, this alignment is not necessarily guaranteed for each question, partly because the WVS is conducted every five years as repeated cross-sectional surveys, with each wave consisting of a different set of respondents, and the Matching-LLM responses are weighted using the responses of WVS6 participants who are demographically most similar to those in WVS7. Table 1 presents the mean absolute deviation (MAD) between LLM or Matching-LLM responses and human responses. Columns (1) and (4) show the differences between LLM responses and human responses for the U.S. and China samples, respectively. Similarly, columns (2) and (5) display the differences between Matching-LLM responses and human responses. The signs of these differences only indicate whether the responses generated by the LLM or Matching-LLM are greater or smaller than the human responses. Last, columns (3) and (6) present the differences between the values in columns (1) and (2), and columns (4) and (5), respectively, to examine whether the Matching-LLM approach outperforms the LLM approach.
Table 1: Matching-LLM outperforms LLM in mean comparison of responses
U.S. China
LLM Matching-LLM LLM Matching-LLM
Survey Question MAD MAD Difference MAD MAD Difference
@@
Male-preferred employment 0.222 0.093 0.129∗∗∗ 0.056 0.010 0.046
Domestic worker priority 0.103 0.059 0.044 0.434 0.295 0.139∗∗∗
Male as breadwinner 0.240 0.148 0.092∗∗∗ 0.190 0.124 0.066∗∗
@@
Trust on family 0.404 0.354 0.050∗∗∗ 0.143 0.114 0.029∗∗∗
Trust on neighbors 0.195 0.166 0.029 0.014 0.015 0.001
Trust on acquaintances 0.178 0.142 0.036∗∗ 0.079 0.019 0.060∗∗∗
Trust on strangers 0.279 0.223 0.056∗∗∗ 0.046 0.012 0.034∗∗
Panel C: Ethical norms and values
Fare evasion acceptance 1.425 0.88 0.545∗∗∗ 0.200 0.404 0.204∗∗∗
Theft acceptance 0.851 0.59 0.261∗∗∗ 0.293 0.060 0.287∗∗∗
Tax evasion acceptance 0.770 0.481 0.289∗∗∗ 0.378 0.505 0.127∗∗∗
Bribery acceptance 0.717 0.42 0.297∗∗∗ 0.512 0.247 0.265∗∗∗
Homosexuality acceptance 1.482 0.873 0.609∗∗∗ 3.331 2.347 0.984∗∗∗
Divorce acceptance 0.301 0.007 0.294∗∗∗ 1.603 1.257 0.346∗∗∗
Parental abuse acceptance 0.926 0.625 0.301∗∗∗ 0.974 0.305 0.669∗∗∗
Note: *, **, and *** represent significance at 10, 5, and 1% levels, respectively.
We found that for all survey questions, the Matching-LLM responses in the U.S. sample were significantly closer to human responses compared to the LLM responses. A similar pattern was observed for the China sample overall, though with some exceptions. For instance, in the cases of “Fare evasion acceptance” and “Tax evasion acceptance,” the Matching-LLM results were significantly further from human responses. Additionally, for questions like “Male-preferred employment” and “Trust on neighbors,” the differences were not statistically significant. Given the five-year interval between the two WVS waves, we believe that the weighted results observed in China are more likely due to greater changes in social norms compared to the U.S. during this period. Overall, the results in Table 1 demonstrate that Matching-LLM significantly improves the accuracy of mean comparisons in responses compared to using LLM alone. This finding lends support to our approach of combining fast thinking (historical factors) with slow thinking (LLM).
U.S.-China differences. Figure 2 shows differences between U.S. and China responses. We observe that the LLM and Matching-LLM replicate most actual U.S.-China differences in human responses, capturing both the sign of the differences and their statistical significance. There are also some exceptions. For instance, LLM-generated U.S.-China differences in “Trust on acquaintances” and “Trust on strangers” diverge in predicted direction. Nevertheless, the Matching-LLM approach effectively corrects this divergence. In terms of common-sense knowledge, LLM responses are stable and unrelated to U.S.-China differences.
Figure 2: U.S.-China differences
Regarding questions related to ethical norms and laws, the LLM inconsistently captures U.S.-China differences, retaining them only for “Homosexuality acceptance”, “Divorce acceptance”, and “Parental abuse acceptance”. The weaker performance in these questions may stem from the LLM’s tendency to align with established laws, which can cause it to diverge from real human responses. However, the Matching-LLM preserves the U.S.-China difference in “Theft acceptance” and, for questions where the LLM simulated an incorrect direction of difference, even without correcting the direction, the Matching-LLM provides differences with smaller absolute values.
@@
Pairwise correlation. Researchers often care about correlations between behaviors and opinions. Due to budget constraints, they often measure multiple preferences in one survey or experiment (e.g., trust and risk-taking), making the consistency of pairwise correlations among these measurements important. Therefore, the LLM should generate responses that not only have means similar to human responses but also preserve their correlations at the individual level.
@@
We assess whether statistically significant pairwise correlations observed between two questions in the LLM-generated responses are likewise present in human responses (Snowberg and Yariv,, 2021). The results, shown in Figure 3, display the coefficients and significance of these pairwise correlations observed among questions on social values and trust. Each circle in the figure represents the correlation between a pair of questions. The horizontal axis indicates the correlation coefficients for these two questions in the human sample, while the vertical axis represents the correlation coefficients for these two questions in the LLM-generated sample (solid circles) or the Matching-LLM sample (hollow circles). Circles where pairwise correlations agree in both sign and significance between LLM-simulated and human samples are labeled as “complete agreement” and shaded red. Circles where the correlation is significant in one sample but insignificant in the other are labeled “partial disagreement” and shaded green. Lastly, circles with significant correlations in both samples but in opposite directions are marked as “complete disagreement” and shaded blue. We do not include common-sense ethical norms and values questions because LLM-simulated responses lack variation.